(D+3)^2y=sinh2x


1
Expert's answer
2021-06-14T17:44:19-0400

The auxiliary equation to the homogenous part of the given differential equation is given as:


"(D+3)^2=0\\"

Thus, we proceed to get the root of the equation:


"(D+3)(D+3)=0\\
D=-3 \ (twice)"

"\therefore C \cdot F .=\left(C_{1}+C_{2} x\right) e^{-3 x}\\
\text{}\\
\operatorname{Sinh} x=\frac{e^{x}-e^{-x}}{2}\\
\text{}\\
\text{Particular Integral}\\
P.I. =\frac{\operatorname{Sinh} 2 x}{(D+3)^{2}}\\
=\frac{1}{2}\left[\frac{e^{2 x}-e^{-2 x}}{(D+3)^{2}}\right]\\
=\frac{1}{2} \frac{e^{2 x}}{(D+3)^{2}}-\frac{1}{2} \frac{e^{-2 x}}{(D+3)^{2}}\\
=\frac{1}{2} \frac{e^{2 x}}{(2+3)^{2}}-\frac{1}{2} \frac{e^{-2 x}}{(-2+3)^{2}}\\
=\frac{e^{2 x}}{50}-\frac{e^{-2 x}}{2}\\
\text{}\\
\text{Hence the complete solution}\\
Y=C . F .+P . I\\
Y=\left(C_{1}+C_{2} x\right) e^{-3 x}+\frac{e^{2 x}}{50}-\frac{e^{-2 x}}{2}\\"